On August 3, Cinco Construction purchased special-purpose equipment at a cost of $9,911,000. The useful life of the equipment was estimated to be eight years, with an estimated residual value of $26,750.

a. Compute the depreciation expense to be recognized each calendar year for financial reporting purposes under the straight-line depreciation method (half-year convention).

b. Compute the depreciation expense to be recognized each calendar year for financial reporting purposes under the 200 percent declining-balance method (half-year convention) with a switch to straight-line when it will maximize depreciation expense.

c. Which of these two depreciation methods (straight-line or double-declining-balance) results in the highest net income for financial reporting purposes during the first two years of the equipment’s use?
Answer #1

a. Depreciation expense per year = [9911000 -26750] / 8

= $1235531

Year Depreciation expense
1 1235531/2 = $617766
2 $1235531
3 $1235531
4 $1235531
5 $1235531
6 $1235531
7 $1235531
8 $1235531
9 1235531/2 = $617766
Total 9884249

b.   200 percent declining-balance rate = 1/8 * 200%

= 25%

Year Book value at beginning [Cost - Depreciation expense per year] Depreciation expense
1 $9,911,000 9911000*25%*1/2 = 1238875
2 8672125 8672125 *25% = 2168031
3 6504094 6504094*25% =1626024
4 4878070 4878070*25% = 1219518
5 3658553 3658553 *25% = 914638
6 2743915 2743915 *25% =685979
7 2057936 2057936*25% = 514484
8 1543452 1543452*25% = 385863
9 1157589 1130839
Total 9884250

3. straight-line method results in the highest net income as its two year depreciation expense is lesser than the double-declining-balance)